Allentown Post is moving to a paywall subscription service rather than a free news website with unlimited access. If subscribers would like to access more than 10 articles per month, they will need to pay a monthly subscription fee of $20. However, if they are also weekly subscribers of the print edition of the newspaper, they receive a 45% discount on the online subscription rate. The monthly rate for the print edition is $30. Based on market research, the Post believes that 25% of who order the print edition will also purchase the online subscription. While there are basically no variable costs to the website version, the print edition does cost $15 per month per subscriber.

c) What is the value of an online-only subscriber presuming s/he keeps the subscription for the full 3 years?
Answer #1

Subscription fee for online -only subscriber =$ 20 per month

Number of months the user is subscribed =3 years =3*12=36 months

Therefore ,

Value of online-only subscriber =Subscription fee/month * Number of months

=20*36

=$ 720